# Hidden long-range correlations in the ion distribution at the graphite / [bmim][NTf2] electrified interface

### v1

@@ III.3 Subtle electrostatic correlations deep into the bulk-like interior of the electrolyte
Acceptance - rejection of the bond interchange is decided by the usual Metropolis rule at fictitious temperature $T$, whose progressive reduction (simulated annealing) drives the system towards the minimum. The scale of he temperature interval and the schedule of the annealing have been decided by trial and error using a short sequence of preliminary computations. Measuring charges in units of $e$ and distances in nm (a unit dipole in these units is equal to 48 Debye units), temperature has been decreased from $T=2$ to $T=0.025$ energy units in 80 runs of $64\ 10^{6}$ MC steps, each consisting of the attempted interchange of bonds between neighbouring dipoles. Since the attempted MC move is a discrete event that cannot be fractioned, the acceptance ratio of the MC cannot be tuned, but it exceeded $0.5$ at high $T$, and decreased to $5\ 10^{-5}$ (one accepted move every 20 thousand moves) at the lowest $T=0.025$.
Figure 9: Dipoles’ self energy as a function of $T$ during the annealing.
The effect of the MC annealing on the definition of dipoles on top of a single configuration of the ions is illustrated in Fig. 8. Arrow represent the strength and spatial location of each dipole in the sample. The choice of the location of each dipole is explained in the following paragraph. The bottom panel shows the final configuration resulting from the simulated annealing. The configuration used as an example has been extracted from the simulation of the capacitor at $T=300$ K, $P=1$ bar, $\sigma=0$. The topmost panel shown a configuration of dipoles selected at random during the simulation at the highest temperature ($T=2$) of the simulated annealing protocol. The color of each arrow is assigned as follows: if the $x$ component of the dipole points to the right, the arrow is painted red, otherwise it is painted gray. The number of red dipoles equals the number of gray dipoles to within $1$ %. From the figure, it is apparent that at the end of the MC procedure, dipoles are of similar size and the sum of their self-energies, although unlikely to be the true absolute minimum, is already low in energy (see the self energy as a function of $T$ during the annealing in Fig. 9). It might be worth emphasising that this procedure is carried out at fixed position of all atoms in the system, only the distribution of fictitious dipole bonds among ions being interchanged.
@@
This allows to compute the density of dipole moment along each direction, which is displayed in Fig. 11 (a) for the sample illustrated in Fig. 8, having $\sigma=0$ surface charge density on the graphite electrodes. It is apparent that the distribution of dipole moments vanishes over most of the sample, with possibly a narrow peak in proximity of the two interfaces.
It can be supposed that a variety of subdivisions into pairs have comparable and relatively low self-energy, all representing the properties of the ideal unique and best pairing choice. However, repeated pairing optimisations, carried out by starting from different randomized starting point had a minimum fraction of equal pairs at the end of the simulated annealing of 66 %, showing that the self-energy basin around the (unknown) ideal minimum is relatively wide, but also well defined, without a plurality of significantly different local minima making ambiguous the pairing of ions into neutral pairs.
@@
The same protocol could be used to pairs ions into dipoles when the capacitor plates carry a non vanishing surface charge density. The resulting configuration of dipoles, however, is less regular than the one obtained at $\sigma=0$, and the MC optimisation is less able to provide a low energy partition of ions in dipoles of nearly homogeneous length. The reason is that the excess of ions of a given $\pm e$ charge screening the $\mp\sigma$ charge on the plate cannot find a neibouring partner in their immediate vicinity, being forced to form dipoles with ions located far into the IL bulk, or even close to the opposite plate, there an excess of counter ions can be found. In fact, at the end of virtually every simulated annealing cycles, the dipoles population included a few long separation (high self-energy) dipoles, unable to escape a local minimum of the total self-energy, and only a sustained effort is able to cure these configurations through very long simulated annealing cycles.
To overcome this problem, the original ions’ population is supplemented by further ions, whose role is to represent the charge on the capacitor plates, providing a pairing opportunity for the counter-ions that populate the first IL layer next to the interface interface. This procedure is better illustrated with an example. Let us consider the same IL sample of $3072$ ion pairs, enclosed in between two graphite electrodes whose interfacial layers (each consisting of $4784$ carbon atoms) now carry $\pm 0.05\ e$ for a total of $\pm 239.2$ e on the two capacitor plates. In the simulation, the left plate is negatively charged, and the rightmost plate is positively charged. Since each carbon atoms carries only $\pm 0.05$ e, the representation of the charge on each electrode cannot be obtained by including all these atoms in the dipole formation protocol. Instead, a number of integer charges corresponding to the total charge on each electrode is introduced. In the present example $239$ e-e charges are located on the leftmost electrode, and 239 +e+e charges are positioned on the rightmost electrode.
@@
After this first $239+239$ pairs are formed, linking ions of opposite sign across the interface, the remaining 2594 ion pairs are given an initial pairing into dipoles precisely as in the $\sigma=0$ case. Then, the MC annealing is carried out following the same schedule as in the neutral case. Bond interchanges concern both the original ions, and the ions added to represent the surface charge.
Following this approach, the optimization of the self energy runs much smoothly and as effectively as in the neutral case, producing a probability distribution of dipole separations $R_{d}$ virtually superimposable to the one from the neutral case. The only difference is a tail at short separations apparent in the $\sigma=2$ e/nm2 plot, due to the tight pairs joining the added surface charges to the counter-ions in the inner screening layer (see Fig. 10). In the annealed configuration, 85 % of the dipoles point towards the right electrode. As apparent from the figure, the dipole density has a net excess of $d_{x}$ dipole moment along the entire capacitor along $x$. The net sum of all dipoles along $x$ has a direction which is such to partly balance the dipole formed by the two graphite surfaces of $\pm\sigma$ charge density. The observation of an excess of dipole moment far from charged interfaces is the major result of the present investigation.
Also in this case, repeating the simulated annealing after randomising the pairing of ions results in a set of ion pairs whose overlap with any previous one is very sizeble. In fact, the overlap seems to be even stronger than in the neutral case, perhaps suggesting that the minimum energy basin is slightly narrower and somewhat simpler for charged interfaces than for the neutral ones.
Figure 12: Dipoles’ configuration at he end of the simulated annealing optimization as a function of $\sigma$. Red dipoles point to the right side of the figure, white dipoles point to the left. The prevailing direction of dipoles in the charged interface case is such to partially compensate the dipole moment of the two charged graphite layers forming the plates of the capacitor.
A quantitative assessment of the dipoles’ alignemenmt along $x$ is given in Fig. 11, showing the average dipole moment along the directions parallel andnormal to the interface. The data reported in this figure refer to which is displayed in Fig. 11 (b) for the sample illustrated in Fig. 12, having $\sigma=2$ surface charge density on the graphite electrodes. It is apparent that the distribution of the $d_{x}$ component of the dipole moment does not vanish over the entire with of the capacitor.
@@ IV Summary and discussion
The interface between ionic liquids and solid electrodes is a fascinating subject that joins the current interest for a vast number of electrochemical applications with concepts and theories dating back to one century ago.korn1 ; korn2 ; korn3 ; bock . This fertile combination, together with the complex structure of ions and the nanostructures texture of ionic liquids, has fostered a variety of results of practical interest, and also some surprises that challenge the accepted picture of the electrified interfaces. Among the aspects that are being challenged is the long held notion that, beyond a narrow interfacial region of microscopic width, a Coulombic fluid in contact with a charged interface is virtually bulk-like. A strong indication in this direction is provided by the recent observationexpLangmuir of the Pockel effect at the interface between the [bmim][NTf2] ionic liquid and two planar solid parallel electrodes connected to a generator mantaining an electrostatic potential energy difference between the two sides of the IL slab. Upon charging the interface, a shift in the Raman frequencies was observed, arising from all Raman-active modes of cations and anions, independently from their distance from the interface.
@@
To investigate this phenomenon, the interface between the same [bmim][NTf2] ionic liquid and a graphite electrode has been simulated by molecular dynamics, using an atomistic semi-empirical force field. An innovative analysis of trajectories allows to join [bmim]+ and [NTf2]- ions into neutral pairs, whose dipoles show an apparent alignement with the external electric field driven by the charge deposited on the graphite surface. This alignment is observed even well inside the IL bulk-like region, where the external field is exactly cancelled (on average) by the screening at the interface by the mobile ions in the IL. Since dipoles do not obey the strict screening conditions of charges, the observation does not violate any accepted theory of electrified interfaces, but opens the way to a whole new class of medium-long range correlations that remain hidden when one considers the system only as a collection of free ions.
@@
Up to now, this observation concerns a geometric property in the distribution of ions throughout the system, that leaves open questions concerning its origin and its significance with respect to the observed effect on the Raman spectra measured on the experimental samples. The main question on its origin concerns the role of the shape anisotropy of the ions in transferring the information on the state of the interface deep into the IL bulk. The main question on the relevance of the geometric property concerns whether this correlation among virtual dipoles can alter the vibrational properties of the ionic liquid. The mechanism joining the geometric correlation with the change of Raman frequences has to respect the fact that the average electric field well inside the ionic liquid vanishes, independently from the charge on the electrode. This fact, however, does not prevent changes in the average square amplitude of the fluctuating electric field inside the ionic liquid, that could indeed give origin to the observed change of vibratioal frequences.
@@
Because of these open questions that are currently being investigated, the present document represents a report on work in progress, that will be followed by a full manuscriprit on the comprehensive investigation of the [bmim][NTf2] / graphite interface and on the hidden correlations in the distribution of ions arising upon chaarging the electrodes.

### v2

@@ III.3 Subtle electrostatic correlations deep into the bulk-like interior of the electrolyte
Acceptance - rejection of the bond interchange is decided by the usual Metropolis rule at fictitious temperature $T$, whose progressive reduction (simulated annealing) drives the system towards the minimum. The scale of he temperature interval and the schedule of the annealing have been decided by trial and error using a short sequence of preliminary computations. For the sake of simplicity, electrostatic and energy units will be expresses in scaled units, in which the unit of length is the nm, the unit of charge is $e$ and the unit of energy is $e^{2}/(1nm)$, corresponding to $0.0529$ Ha, or 1.44 eV, or $139$ kJ/mol. In these units, a unit dipole is $48$ Debye, a unit electric field is the one producing a unit energy change on a single positive $e$ charge upon a 1 nm displacement, etc. Also, measuring temperature in energy units, $T$ has been decreased from $T=2$ to $T=0.025$ energy units in 80 runs of $64\ 10^{6}$ MC steps, each consisting of the attempted interchange of bonds between neighbouring dipoles. Since the attempted MC move is a discrete event that cannot be fractioned, the acceptance ratio of the MC cannot be tuned, but it exceeded $0.5$ at high $T$, and decreased to $5\ 10^{-5}$ (one accepted move every 20 thousand moves) at the lowest $T=0.025$.
Figure 9: Dipoles’ self-energy $u_{s}$ as a function of $T$ during the annealing. Uncharged capacitor at $T=300$ K. The scaled units for $u_{s}$ and $T$ are defined in the text. The $u_{s}$ values refer to the self energy of all $3072$ ion pairs.
The effect of the MC annealing on the definition of dipoles on top of a single configuration of the ions is illustrated in Fig. 8. Arrow represent the strength and spatial location of each dipole in the sample. The choice of the location of each dipole is explained in the following paragraph. The bottom panel shows the final configuration resulting from the simulated annealing. The configuration used as an example has been extracted from the simulation of the capacitor at $T=300$ K, $P=1$ bar, $\sigma=0$. The topmost panel shown a configuration of dipoles selected at random during the simulation at the highest temperature ($T=2$) of the simulated annealing protocol. The color of each arrow is assigned as follows: if the $x$ component of the dipole points to the right, the arrow is painted red, otherwise it is painted gray. The number of red dipoles equals the number of gray dipoles to within $1$ %. From the figure, it is apparent that at the end of the MC procedure, dipoles are of similar size and the sum of their self-energies, although unlikely to be the true absolute minimum, is already low in energy (see the self-energy as a function of $T$ during the annealing in Fig. 9). It might be worth emphasising that this procedure is carried out on a single configuration of the sample, i.e., at fixed position of all atoms in the system, only the distribution of fictitious dipole bonds among ions being interchanged.
@@
It can be supposed that a variety of subdivisions into pairs have comparable and relatively low self-energy, all representing the properties of the ideal unique and best pairing choice. However, repeated pairing optimisations, carried out by starting from different randomized starting point had a minimum fraction of equal pairs at the end of the simulated annealing of 66 %, showing that the self-energy basin around the (unknown) ideal minimum is relatively wide, but still well defined, without a plurality of significantly different local minima making ambiguous the pairing of ions into neutral pairs.
Figure 12: Self-energy $u_{s}$ of the optimal ion pairing determined for the MD configuration at time $\tau_{0}$, ri-evaluated on the MD configuration at time $\tau>\tau_{0}$. In other terms, a list of ions pairs minimising the self-energy $u_{s}$ is determined at time $\tau_{0}$. Then, the self-energy expression for the same ion pairs is ri-evaluated using the ions’ coordinates extracted from the MD trajectory at time $\tau$. The $u_{s}$ value refers to the total population of $3072$ ion pairs.
Since the grouping of ions into dimers is the result of a probabilistic algorithm that runs independently on different configurations, it is difficult to estimate an accurate value for the dipoles’ lifetime. However, it has been verified that the ion pairing resulting from optimisation protocol on the MD configuration at time $\tau_{0}$ remains a low self-energy pairing for ions’ configurations extracted at later times $\tau>\tau_{0}$ for $\tau-\tau_{0}$ of the order of tens of ns, as can be seen in Fig. 12. This observation suggests a lifetime of a similar duration for the dipoles.
@@
This allows to compute the density of dipole moment along each direction, which is displayed in Fig. 11 (a) for the sample illustrated in Fig. 8, having $\sigma=0$ surface charge density on the graphite electrodes. It is apparent that the distribution of dipole moments vanishes over most of the sample, with possibly narrow peaks in proximity of the two interfaces.
@@
The same protocol could be used to pairs ions into dipoles when the capacitor plates carry a non vanishing surface charge density. The resulting configuration of dipoles, however, is less regular than the one obtained at $\sigma=0$, and the MC optimisation is less able to provide a low energy partition of ions in dipoles of nearly homogeneous length. The reason is that the excess of ions of a given $\pm e$ charge screening the $\mp\sigma$ charge on the plate cannot find a neighbouring partner in their immediate vicinity, being forced to form dipoles with ions located far into the IL bulk, or even close to the opposite plate, there an excess of counter ions can be found. In fact, at the end of virtually every simulated annealing cycles, the dipoles population included a few long separation (high self-energy) dipoles, unable to escape a local minimum of the total self-energy, and only a sustained effort is able to cure these configurations through very long simulated annealing cycles.
To overcome this problem, the original ions’ population is supplemented by further ions, whose role is to represent the charge on the capacitor plates, providing a pairing opportunity for the counter-ions that populate the first IL layer next to the interface. This procedure is better illustrated with an example. Let us consider the same IL sample of $3072$ ion pairs, enclosed in between two graphite electrodes whose interfacial layers, each consisting of $4784$ carbon atoms, now carry $\pm 0.05\ e$ per C atom, for a total of $\pm 239.2$ e on the two capacitor plates. In the simulation, the left plate is negatively charged, and the right plate is positively charged. Since each carbon atoms carries only $\pm 0.05$ e, the representation of the charge on each electrode cannot be obtained by including all these atoms in the dipole formation protocol. Instead, a number of integer charges corresponding to the total charge on each electrode is introduced. In the present example $239$ e-e charges are located on the left electrode, and 239 +e+e charges are positioned on the right electrode.
@@
After this first $239+239$ ion pairs are formed, linking ions of opposite sign across the interface, the remaining 2594 ion pairs are given an initial pairing into dipoles precisely as in the $\sigma=0$ case. Then, the MC annealing is carried out following the same schedule as in the neutral case. Bond interchanges concern both the original ions, and the ions added to represent the surface charge.
Following this approach, the optimization of the self-energy runs as smoothly and as effectively as in the neutral case, producing a probability distribution of dipole separations $R_{d}$ virtually superimposable to the one from the neutral case. The only noticeable difference is a tail at short separations apparent in the $\sigma=2$ e/nm2 plot, due to the tight pairs joining the added surface charges to the counter-ions in the inner screening layer (see Fig. 10). In the annealed configuration, 85 % of the dipoles point towards the right electrode. As apparent from the figure, the dipole density has a net excess of $d_{x}$ dipole moment along the entire capacitor along $x$. The net sum of all dipoles along $x$ has a direction which is such to partly balance the dipole formed by the two graphite surfaces of $\pm\sigma$ charge density. The observation of an excess of dipole moment far from charged interfaces is the major result of the present investigation.
Also in this case, repeating the simulated annealing after randomising the pairing of ions results in a set of ion pairs whose overlap with any previous one is very sizeable. In fact, the overlap seems to be even stronger than in the neutral case, perhaps suggesting that the minimum energy basin is slightly narrower and somewhat simpler for charged interfaces than for the neutral ones.
Figure 13: Dipoles’ configuration at he end of the simulated annealing optimization as a function of surface charge density $\sigma$. Red dipoles point to the right side of the figure, white dipoles point to the left. The prevailing direction of dipoles in the charged interface case is such to partially compensate the dipole moment of the two charged graphite layers forming the plates of the capacitor.
A quantitative assessment of the dipoles’ alignment along $x$ is given in Fig. 11, showing the average dipole moment along the directions parallel and normal to the interface. The data reported in this figure refer to which is displayed in Fig. 11 (b) for the sample illustrated in Fig. 13, having $\sigma=2$ surface charge density on the graphite electrodes. It is apparent that the distribution of the $d_{x}$ component of the dipole moment does not vanish over the entire with of the capacitor.
@@ IV Summary and discussion
The interface between ionic liquids and solid electrodes is a fascinating subject that joins the current interest for a vast number of electrochemical applications with concepts and theories dating back to one century ago.korn1 ; korn2 ; korn3 ; bock . This fertile combination, together with the complex structure of ions and the nanostructures texture of ionic liquids, has fostered a variety of results of practical interest, and also some surprises that challenge the accepted picture of the electrified interfaces. Among the aspects that are being challenged is the long held notion that, beyond a narrow interfacial region of microscopic width, a Coulombic fluid in contact with a charged interface is virtually bulk-like. A strong indication in this direction is provided by the recent observationexpLangmuir of the Pockel effect at the interface between the [bmim][NTf2] ionic liquid and two planar solid parallel electrodes connected to a generator maintaining an electrostatic potential energy difference between the two sides of the IL slab. Upon charging the interface, a change in the Raman intensity was observed, arising from all Raman-active modes of cations and anions, independently from their distance from the interface.
@@
To investigate this phenomenon, the interface between the same [bmim][NTf2] ionic liquid and a graphite electrode has been simulated by molecular dynamics, using an atomistic semi-empirical force field. An innovative analysis of trajectories allows to join [bmim]+ and [NTf2]- ions into neutral pairs, whose dipoles show an apparent alignment with the external electric field driven by the charge deposited on the graphite surface. This alignment is observed even well inside the IL bulk-like region, where the external field is exactly cancelled (on average) by the screening at the interface by the mobile ions in the IL. Since dipoles do not obey the strict screening conditions of charges, the observation does not violate any accepted theory of electrified interfaces, but opens the way to a whole new class of medium-long range correlations that remain hidden when one considers the system only as a collection of free ions.
@@
Up to now, this observation concerns a geometric property in the distribution of ions throughout the system, that leaves open questions concerning its origin and its significance with respect to the observed effect on the Raman spectra measured on the experimental samples. The main question on its origin concerns the role of the shape anisotropy of the ions in transferring the information on the state of the interface deep into the IL bulk. The main question on the relevance of the geometric property concerns whether this correlation among virtual dipoles can alter the vibrational properties of the ionic liquid. The mechanism joining the geometric correlation with the change of Raman spectroscopic properties has to respect the fact that the average electric field well inside the ionic liquid vanishes, independently from the charge on the electrode. This fact, however, does not prevent changes in the average square amplitude of the fluctuating electric field inside the ionic liquid, that could indeed give origin to the observed change of vibrational frequencies.
@@
Because of these open questions that are currently being investigated, the present document represents a report on work in progress, that will be followed by a full manuscriprit on the comprehensive investigation of the [bmim][NTf2] / graphite interface and on the hidden correlations in the distribution of ions arising upon charging the electrodes.
Further considerations that will be explored during later stages of our investigations are as follows. Long range correlations imply that the two electrodes in simulated planar capacitors are likely to interact up to the largest sizes that are practically achievable right now, affecting any estimate of properties that depend on the spatial distribution and orientation of the ions. This problem enhances the appeal of single-electrode schemes like the one introduced in Ref. eletransf, , in which a unique spherical electrode surrounds the pool of electrolyte, whose thermodynamic state is controlled by adopting the grand canonical simulation approach.
The observation of a peculiar geometric property that conveys information on polarisation from the interface to deep into the bulk electrolyte is reminiscent of the role of the Berry phase in ab-initio investigations of polarisation in ferroelectric crystals.raffaele Although the Berry phase is, to our knowledge, an exclusively quantum mechanical feature, it might be worth exploring the analogy with the long range correlations highlighted in the present study.
ACKNOWLEDGMENTS: N.C.F.-M. gratefully acknowledges the ECHELON Project: P2019-02-003 from the Carl Zeiss Foundation for financial support. D.V.-D., and N.C.F.-M. acknowledge the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) for funding through the research group FOR 2982–UNODE, Project number 413163866. R.C.-H. acknowledges funding from SFB-TRR146 of the German Research Foundation (DFG)–Project No. 233630050 and the European Union funding through the Twinning project FORGREENSOFT (grant no. 101078989 under HORIZON-WIDERA-2021-ACCESS-03). D.V.-D., R.C.-H., and N.C.F.-M. thank Kurt Kremer and Friederike Schmid for useful discussions.